# A FLASH on Blazars

### v2

@@ 6.1 Summary
This study explores the radio properties of 4FGL blazars detected in the FLASH continuum survey, focusing on their spectral characteristics and multi-wavelength correlations. The main findings are summarized as follows:
@@
  • Extended Emission: At low frequencies (e.g., below 100 MHz), the emission is characterised by a spectral index of $\sim$-∼ -0.54, consistent with the presence of an additional steep-spectrum emission, previously missed due to the lack of high-resolution and deep low frequency data .
  • Radio-gamma correlation: A strong correlation between radio and gamma-ray luminosities was identified, particularly at 855 MHz, emphasizing the dominant contribution of relativistic jets to the observed multi-wavelength emission.
  • Radio-X-ray correlation: The X-ray luminosity (0.2–2.0 keV) is tightly correlated with radio luminosity for FSRQs, reflecting their accretion-driven nature, while BL-Lacs show a more scattered relationship, consistent with their jet-driven dynamics.
  • Blazar Subclass Differences: FSRQs display more efficient high-energy pumping and dominate the bright end of the radio and gamma luminosity spectrum, while BL-Lacs exhibit greater variability in spectral behavior, reflecting their evolutionary differences.
@@ 6.2 Future perspectives
Building on these findings, future studies will aim to deepen our understanding of the physical and environmental properties of blazars. Firstly, FLASH is an ongoing survey; therefore, we will expand the blazar sample to include the full 5-year dataset upon completion. Further, a key next step will be to analyse the spectroscopic properties of these objects using FLASH, XMM-Newton, and optical lines data (e.g., broad-line region emission). This will help us provide detailed insights into the physics of blazars and their environments, offering clues about the retriggering mechanism and the dual nature of their radio emission. Finally, leveraging instruments like the SKAO and CTA will allow for broader frequency coverage and improved sensitivity, enabling the detection of faint sources and the study of synchrotron emission at unprecedented depth.
@@ Acknowledgements.
This work was carried out in part during an Erasmus+ exchange program at CSIRO ATNF in Marsfield, Sydney. MB gratefully acknowledges the support provided by Erasmus+ and the CSIRO ATNF for facilitating this research opportunity. This scientific work uses data obtained from Inyarrimanha Ilgari Bundara, the CSIRO Murchison Radio-astronomy Observatory. We acknowledge the Wajarri Yamaji People as the Traditional Owners and native title holders of the Observatory site. CSIRO’s ASKAP radio telescope is part of the Australia Telescope National Facility (https://ror.org/05qajvd42). Operation of ASKAP is funded by the Australian Government with support from the National Collaborative Research Infrastructure Strategy. ASKAP uses the resources of the Pawsey Supercomputing Research Centre. Establishment of ASKAP, Inyarrimanha Ilgari Bundara, the CSIRO Murchison Radio-astronomy Observatory and the Pawsey Supercomputing Research Centre are initiatives of the Australian Government, with support from the Government of Western Australia and the Science and Industry Endowment Fund. This work was partially funded from the projects: INAF GO-GTO Normal 2023 funding scheme with the project ”Serendipitous H-ATLAS-fields Observations of Radio Extragalactic Sources (SHORES)”. MVZ acknowledges partial financial support from the ”Fondazione CR Firenze” with the program ”Ricercatori a Firenze 2023” The colour palettes used in this paper were directly extracted from Claude Monet’s paintings Water Lilies and San Giorgio Maggiore at Dusk.
See pages -12 of all_seds.pdf See pages 13- of all_seds.pdf

### v3

@@ 6.1 Summary
In this study, we have explored the radio properties of 4FGL blazars detected in the FLASH continuum survey while focusing on their spectral characteristics and multi-wavelength correlations. The main findings are summarised as follows:
@@
  • Extended emission: At low frequencies (e.g. below 100 MHz), the emission is characterised by a spectral index of $\sim$-0.54, which is consistent with the presence of an additional steep-spectrum emission that was previously missed due to a lack of high-resolution and deep low-frequency data.
  • Radio-gamma correlation: A strong correlation between radio and gamma-ray luminosities was identified, particularly at 855 MHz, emphasising the dominant contribution of relativistic jets to the observed multi-wavelength emission.
  • Radio–X-ray correlation: The X-ray luminosity (0.2–2.0 keV) is tightly correlated with radio luminosity for FSRQs, reflecting their accretion-driven nature, while BL Lacs show a more scattered relationship, which is consistent with their jet-driven dynamics.
  • Blazar subclass differences: FSRQs display more efficient high-energy pumping and dominate the bright end of the radio and gamma-ray luminosity spectrum, while BL Lacs exhibit greater variability in spectral behaviour, reflecting their evolutionary differences.
@@ 6.2 Future perspectives
Building on these findings, future studies will be aimed at deepening our understanding of the physical and environmental properties of blazars. As FLASH is an ongoing survey, we will expand the blazar sample to include the full five-year dataset upon its completion. Further, a key next step will be to analyse the spectroscopic properties of these objects using FLASH, XMM-Newton, and optical lines data (e.g. broad-line region emission). This will help us provide detailed insights into the physics of blazars and their environments, and offer clues about the re-triggering mechanism and the dual nature of their radio emission. Finally, leveraging instruments such as the SKA Observatory and Cherenkov Telescope Array will allow for broader frequency coverage and improved sensitivity and thus enable the detection of faint sources and the study of synchrotron emission at unprecedented depth.
7 Data availability
Tables listing the properties of the sources are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/aa54523-25. .
@@ Acknowledgements.
This work was carried out in part during an Erasmus+ exchange program at CSIRO ATNF in Marsfield, Sydney. MB gratefully acknowledges the support provided by Erasmus+ and the CSIRO ATNF for facilitating this research opportunity. This scientific work uses data obtained from Inyarrimanha Ilgari Bundara, the CSIRO Murchison Radio-astronomy Observatory. We acknowledge the Wajarri Yamaji People as the Traditional Owners and native title holders of the Observatory site. CSIRO’s ASKAP radio telescope is part of the Australia Telescope National Facility (https://ror.org/05qajvd42). Operation of ASKAP is funded by the Australian Government with support from the National Collaborative Research Infrastructure Strategy. ASKAP uses the resources of the Pawsey Supercomputing Research Centre. Establishment of ASKAP, Inyarrimanha Ilgari Bundara, the CSIRO Murchison Radio-astronomy Observatory and the Pawsey Supercomputing Research Centre are initiatives of the Australian Government, with support from the Government of Western Australia and the Science and Industry Endowment Fund. This work was partially funded from the projects: INAF GO-GTO Normal 2023 funding scheme with the project ’Serendipitous H-ATLAS-fields Observations of Radio Extragalactic Sources (SHORES)’. MVZ acknowledges partial financial support from the ’Fondazione CR Firenze’ with the program ’Ricercatori a Firenze 2023’. The colour palettes used in this paper were directly extracted from Claude Monet’s paintings Water Lilies and San Giorgio Maggiore at Dusk.